A small electronics company assembles power supplies for use in computers, printers and FAX machines. They want to set up an assembly line to make their operation more efficient and plan to use line balancing with the incremental utilization heuristic. The demand for power supplies is 200 per hour, and the productive time per hour is 52 minutes.

Task Tasks that must immediately preceede time to perform task

A

-

0.56

B A 0.34
C A 0.11
D A 0.09
E B,C 0.43
F D 0.22
G E 0.62
H F,G 0.35
I H 0.21
J H 0.57
K I,J 0.33

A. What is the cycle time?

B. What is the theoretical number of workstations?

C. What tasks go into the first workstation?

D. How many workstations are needed for the entire line?

E. What is the utilization of the production line?
Answer #1

1.

Cycle time = 52 min per hr / 200 per hr

= 0.26 minutes / unit

2.

Sum of task times = 3.83 minutes.

Theoritical minimum number of workstations

= 3.83 *200 /52

= 14.73

3.

A,B and C goes into first workstation.

4.

Workstations needed = 6

5.

Utilization = 14.73 / 6 = 0.921